OpenStudy (anonymous):

everyday number of traffic accidents has a distribution x 0,1,2 and more than two having P(x) 0.6,0.2,0.2,0 respectively independent of the other days. what is the probability that there are more accidents on friday than on thursday ?

OpenStudy (anonymous):

I don't know... :/ sorrryyy probably more accidents of friday bc of the drunk drivers though lols :p

OpenStudy (anonymous):

ok np... thanks anyway :)

OpenStudy (zarkon):

I get .28

OpenStudy (zarkon):

One way to do this...let X=accidents on friday and Y=accidents on friday you want P(X>Y) =P(X=1,Y=0 or X=2,Y=0 or X=2,Y=1)

OpenStudy (zarkon):

you can also use symmetry

OpenStudy (anonymous):

thanks zarkon

OpenStudy (anonymous):

@Zarkon Can you go through the question with Annas rather than just an outline as he is still confused.

OpenStudy (zarkon):

Using symmetry... by symmetry \(P(X>Y)=P(X<Y)\) using the fact that \[1=P(X>Y \text{ or }X=Y \text{ or }X<Y)=P(X>Y)+P(X=Y)+P(X<Y)\] \[=2P(X>Y)+P(X=Y)\] solving for \(P(X>Y)\) we get \[P(X>Y)=\frac{1-P(X=Y)}{2}\] and \[P(X=Y)=P(X=0,Y=0 \text{ or }X=1,Y=1 \text{ or }X=2,Y=2)\] \[=P(X=0,Y=0)+P(X=1,Y=1)+P(X=2,Y=2)\] \[=.6^2+.2^2+.2^2=.44\] so \[P(X>Y)=\frac{1-P(X=Y)}{2}=\frac{1-.44}{2}=\frac{.56}{2}=.28\]